1) Katie's Boat Rental has a marginal profit function of ??(?) = 20? − 50. Yesterday Katie made a profit of $400 on 10 boat rentals. What is the profit function, ?(?), for Katie's Boat Rental?

2) John's truck Repair has a channel of repair videos. The total number of subscribers per week is represented by the function ?(?) = 60√? , where t is weeks after he started the channel.

What is the average number of subscribers on his channel between weeks 4 and 9 after starting the channel?
Answer #1

The marginal profit is: M. P(x) = 200-50 Now, the total profit is given as: P(x) = rex dx +C P(x) = f (20x-50) dx +C P(x) = 2